joint probability density function over the range

Determine the value of c that makes the function f(x,y) = cxy a joint probability density function over the range 0<x<3 and 0<y<x.
Answer #1

this needsto be solved for c. ajoint probability must add up to 1.

=(c/2)

=(c/2)34/4 =1 or

c=8/34 =8/81

answered by: starzy